OpenStudy (anikay):

I need some help with calculus, implicit differentiation. 8(x^2)y+7x(y^2)=-4

OpenStudy (anikay):

\[16xy+8x^2y' + 7y^2+ 14xyy'=0\] is this right so far?

OpenStudy (bibby):

That seems right. Then isolate the dy/dx or y' and solve

OpenStudy (anikay):

mmk, thank you. I've done this in high school I just forgot a lot before college :P sometimes I make stupid mistakes.

OpenStudy (anikay):

\[y'(8x^2+14xy)=-16xy-7y^2\]

OpenStudy (anikay):

\[\frac{ -16xy-7y^2 }{ 8x^2+14xy }=y'\] and that'd be right....

OpenStudy (anikay):

I just need to do my work on this equation thingy, when I write it out I get confused....typing is so much clearer. Curses of a technology generation. well thanks guys :)

OpenStudy (bibby):

You might want to look into wolfram for verification btw. http://www.wolframalpha.com/input/?i=differentiate+8%28x%5E2%29y%2B7x%28y%5E2%29%3D-4 although the negatives seem off. I'll try and figure out why

OpenStudy (anikay):

I have the answer, it matches mine perfectly. I just struggled getting there.

OpenStudy (bibby):

I mean in the future. not just for this example.

OpenStudy (bibby):

never mind, the negatives matched up this time.